The following items are taken from the year-end general ledger accounts of RAVENNA, Inc. The bonds were issued by First Medici and carried a 10.0% coupon, payable annually on Mar 31. The lending indenture includes an acceleration clause triggered wen the year-end current ratio is less than 2.0. The patent was acquired in the current year and there were $150 in writeoffs in November. (CPA adapted)

Bonds Payable (Principal due 6/30/2030) 5,000

Accounts Payable 1,600

AFDA 110

Accrued Expenses 800

SGA Expenses 1,810

Accounts Receivable 2,200

Inventory 3,200

Patent 4,000

Required: Based on the amounts that RAVENNA will report, what is its current ratio?
Answer #1

Current assets = ( Accounts receivable - AFDA ) + Inventory

= (2,200-110)+3,200

= $5,290

Accrued interest = Bonds x Interest rate x time period

= 5,000 x 10% x 9/12

= $375

Current liabilities = Accounts payable + Accrued expenses + Accrued interest

= 1,600+800+375

= 2,775

Current ratio = Current assets / Current liabilities

= 5,290/2,775

= 1.91

Other information given in the question is not relevant for calculation current ratio and hence ignored.

Kindly comment if you need further assistance. Thanks‼!